Circular No. 7930

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002dv IN UGC 11486 B. Beutler and W. D. Li, University of California at Berkeley, report

the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered KAIT images taken on July 1.4 (mag ∼ 18.8) and 2.3 UT (mag ∼ 18.7). SN 2002dv is located at α = 19^{h}49^{m}48^{s}.72, δ = +50^{o}41′57′′.3 (equinox 2000.0), which is 0′′.5 west and 11′′.1 north of the nucleus of UGC 11486. KAIT images taken on June 17.4 (limiting mag ∼ 19.5) and 25.4 (limiting mag ∼ 18.5) showed nothing at this position.

SUPERNOVA 2002ap IN M74 Y. Yoshii and colleagues (cf. IAUC 7826) report the following addi-

tional magnitudes (uncertainties 0.05 mag for the optical BV RI bands in the Johnson-Cousins system and 0.1 mag for the near-infrared JHK bands) obtained with the MAGNUM 2-m telescope, showing that SN 2002ap has faded by ∼ 3 mag from peak brightness: June 12.612 UT, V = 16.14; 12.616, K = 14.74; 13.607, H = 14.98; 13.610, B = 16.73; 14.596, J = 15.67; 14.597, R = 15.43; 18.606, H = 15.02; 18.608, I = 15.13.

COMETS C/2002 L7, C/2002 L8, C/2002 M1, C/2002 M2 (SOHO) Further to IAUC 7922, D. Hammer reports his measurements for ad-

ditional Kreutz comets found on SOHO website C2 images by D. Evans (C/2002 L7), M. Oates (C/2002 L8), and R. Kracht; C/2002 M2 was also visible in the C3 coronagraph. Reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 L7 June 13.746  5^{h}25.^{m}0  +21^{o}29′  2002-M22
C/2002 L8  15.295  5 30.8  +21 28  2002-M23
C/2002 M1  20.663  5 51.5  +21 43  2002-M43
C/2002 M2  17.500  5 37.6  +19 27  2002-M47

COMET 7P/PONS-WINNECKE Total visual magnitude estimates: May 8.44 UT, 13.3 (A. Hale, Cloud-

croft, NM, 0.41-m reflector); 13.45, 13.1 (Hale); 21.43, 12.9 (Hale); June 8.42, 12.0 (Hale); 12.60, 11.7 (D. A. J. Seargent, Cowra, N.S.W., 0.25-m reflector); 20.86, 12.2 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); 21.86, 12.4 (Pearce); 22.90, 12.2 (Pearce).

2002 July 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green